OpenStudy (anonymous):

A cyclist rides her bike at a rate of 8 meters per second. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 3 hours? Do not round your answers.

OpenStudy (anonymous):

Convert the meters into kilometers, then the seconds into minutes, then hours. Therefore: \[\frac {8 meters}{1 second} \times \frac {60 seconds}{1 minute} \times \frac{60 minutes}{1 hour} \times \frac{1 kilometer}{1000 meters}\] Or, easily put: \[\frac{8 \times 60 \times 60}{1000}\] To find the second part, just multiply by the number of hours, or 3.

OpenStudy (anonymous):

so i mulitply that?